# Bivariate phase-type distributions for experience rating in disability insurance

### v2

@@ 4.3. Predictive performance
Figure 4.11 illustrates the loglikelihoods obtained in each of the three cases of interest across all iterations of the simulation study. The visual representation enables a more nuanced understanding of the models’ behavior. More specifically, and consistent with our earlier observations, in Case A, the simple and phase-type models excel in all iterations. In contrast, the hierarchical model falls behind, which is expected given that this model can only describe positive correlations in the random effects. Nevertheless, Case B showcases the efficiency of the hierarchical model in capturing positive correlation, with the phase-type model also demonstrating strong performance and even outperforming the hierarchical model in several iterations. Finally, in Case C, the phase-type model’s ability to handle negative correlation sets it apart from the others, leading to the highest loglikelihoods for all iterations.
@@
To evaluate the predictive performance of the models, we carry out premium calculations for a disability rate of one monteary unit until age $67$ under the equivalence principle using both the true parameters and the various estimates. A fourth order Runge-Kutta method with a step length of one month is employed. Premiums are first calculated for integer initial ages, and then linear interpolation is used to determine the premium for non-integer initial ages. Throughout, we use $r\equiv 0.01$, corresponding to a continuously compounded interest rate of $1\%$. The comparative analysis is summarized in Table 4.3, which presents both the mean absolute errors (MAE’s) and the root mean square errors (RMSE’s) for each model under consideration and aggregated across groups and insured.
The findings mirror our previous insights, reinforcing the supremacy of the shrinkage models over the standard and fixed effect approaches. Particularly noteworthy is the phase-type model’s dominance, exhibiting the most favorable performance across all cases (A, B, and C). Intriguingly, despite its earlier promising indications, the hierarchical model lags behind in all instances, even in the case of positive correlation, where a larger loglikelihood had otherwise been observed.
@@
Hierarchical $5.787$ $4.712$ $7.885$ $3.381$ $3.112$ $5.090$
Phase-type $5.512$ $4.511$ $6.008$ $3.114$ $2.975$ $3.517$
Table 4.3. RMSE’s and MAE’s for the total equivalence premium (in hundredths).

### v3

@@ 4.3. Predictive performance
Figure 4.11 illustrates the loglikelihoods obtained in each of the three cases of interest across all iterations of the simulation study. The visual representation enables a more nuanced understanding of the models’ behavior. More specifically, and consistent with our earlier observations, in Case A, the simple and phase-type models excel in all iterations. In contrast, the hierarchical model falls behind, which is expected given that this model can only describe positive correlations in the random effects. Nevertheless, Case B showcases the efficiency of the hierarchical model in capturing positive correlation, with the phase-type model also demonstrating strong performance and even outperforming the hierarchical model in several iterations. Finally, in Case C, the phase-type model’s ability to handle negative correlation sets it apart from the simple model, leading to the highest loglikelihoods for all iterations.
@@
To evaluate the predictive performance of the models, we carry out premium calculations for a disability rate of one monetary unit until age $67$ under the equivalence principle using both the true parameters and the various estimates. A fourth order Runge-Kutta method with a step length of one month is employed. Premiums are first calculated for integer initial ages, and then linear interpolation is used to determine the premium for non-integer initial ages. Throughout, we use $r\equiv 0.01$, corresponding to a continuously compounded interest rate of $1\%$. The comparative analysis is summarized in Table 4.4, which presents both the mean absolute errors (MAE’s) and the root mean square errors (RMSE’s) for each model under consideration and aggregated across groups and insured.
@@
Hierarchical $5.528$ $4.456$ $3.115$ $2.940$
Phase-type $5.512$ $4.509$ $6.006$ $3.114$ $2.974$ $3.517$
Table $4.$ RMSE’s and MAE’s for the total equivalence premium (in hundredths).
The findings mirror our previous insights, reinforcing the supremacy of the shrinkage models over the standard and fixed effect approaches. The phase-type model exhibits the most favorable performance in Cases A and C, while in Case B the hierarchical model takes the lead, with the phase-type model following closely. Overall, this underlines the broader flexibility of the phase-type specification.
Finally, it is essential to note that the hierarchical model’s assumption of equal marginal variances is quite limiting. For example, when considering a scenario where $\Theta_{ai}$ follows a $\text{Gamma}(10/6,10/6)$-distribution and $\Theta_{ia}$ follows a $\text{Gamma}(10/3,10/3)$-distribution, coupled by the same Clayton copula with Kendall’s $\rho_{\tau}=0.5$, even the simple model can outperform the hierarchical model, with an improvement (reduction) in RMSE of about $0.03$. This showcases that, beyond capturing dependence, accounting for differences in marginal variances is also crucial for predictive performance.